OpenStudy (anonymous):

Solve on the interval [0,2pi): 4 sec x + 6 = -2 A. 2pi/3,4pi/3 B. pi/6,5pi/6 C. pi/3,5pi/3 D. 7pi/6,11pi/6

OpenStudy (anonymous):

@RaphaelFilgueiras Why?

OpenStudy (anonymous):

cos(x)=-1/2

OpenStudy (anonymous):

sorry i forget minus

OpenStudy (anonymous):

\[4 \sec x + 6 = -2\]\[4 \sec x = -8\]\[\sec x = -2\]\[\frac{1}{\cos x} = -2\]\[\cos x = - \frac{1}{2}\]\[x = \frac{2 \pi}{3}, \frac{4 \pi}{3}\]

OpenStudy (anonymous):

pi-pi/3,pi+pi/3